OpenStudy (anonymous):

The slope of the line below is -2. Which of the following is the point-slope form of the line? (4,-8)

OpenStudy (anonymous):

is it y - 8 = -2(x + 4) orr....

OpenStudy (rajee_sam):

Point Slope form is written as (y - y1) = m ( x-x1) m = slope and (x1, y1) is the point given to you. So the slope here is -2 and point is (4,-8) Now ( y - y1) will be (y-(-8)) = ?

OpenStudy (anonymous):

Thanks(:

OpenStudy (rajee_sam):

Similarly (x-x1) = (x - 4)

OpenStudy (rajee_sam):

So you got it?

OpenStudy (anonymous):

I think so is it y + 8 = -2(x - 4)?

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

Thanks again:D